Solve the following differential equations

x2d2y/dx2 − xdy/dx − 3y = lnx/ x , x > 0. show that the answer is y = A/x + Bx3 − lnx/ 6x (2lnx + 1)

x d2y/dx2 − dy/dx + 2y/x = (ln x)2 .show that the answer is y = x { A cos(ln x) + B sin(ln x) + (lnx) 2 − 2 }
Answer #1

Solutions 2² dy -ndy -3y= lan noo da? n 27 day lnu ,270 da? 2 This is a day 3y= a evler Cauchy equation which is typically sofor particular integral Y,=e-t x stet Izre3t Walet est = Yezt neot 3e3t Now 92/4, X Yp =-9. y ₂ x dt W et tet dt 4 ezt ter desolution =(nn)? ad y como o The 22 dy ndy +2y= da2 allax)² due du This is a ever couchy equation which is typically solved byfor particular integgel Yp = (At+84 +chet 4$=(2At + 8) et + (1+2+8++Cjet Yp - (za) et + (zat 48) et + (pat + B )et + (At?487+